OpenStudy (anonymous):

a container is 1/3 full of water. if 20 litres is added, then it is 3/4 full. how many litres does the container hold wen full

OpenStudy (anonymous):

48 L

OpenStudy (anonymous):

IS IT CORRECT:-)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no just asking how did you do it??

OpenStudy (anonymous):

l/3 = x 3l/4 = x+20

OpenStudy (anonymous):

1/3x+20=3/4x let x be volume in litres

OpenStudy (anonymous):

you guys are really smart :) thankyou very much

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

mine 2

OpenStudy (anonymous):

wat grade r u guys in

OpenStudy (anonymous):

just passd 12

OpenStudy (anonymous):

weve both passed 12th

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

haha i'm only in grade 6

OpenStudy (anonymous):

ohhh my bro is........:-))))

OpenStudy (anonymous):

my sis is in 7

OpenStudy (anonymous):

r u frm china

OpenStudy (anonymous):

yr name suggests

OpenStudy (anonymous):

my dads name

OpenStudy (anonymous):

and i'm not from china

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

i'm in austrailia

OpenStudy (anonymous):

don't worry

OpenStudy (anonymous):

so ur of oriental descent?

OpenStudy (anonymous):

thank you again i have to go to continue my hw now i might ask u again wen i'm stuck in another question him1618 and sweetznchillz

OpenStudy (anonymous):

jst post it...maybe well b there